Suppose a car rental company charges $110 for the first day and $60 for each additional or partial day. Let S(x) represent the cost of renting a car for x days. Find the value of S(4.5).

Answer by Boreal(15235)

The function is S(x)=110+60(x-1)=50+60x, for x=number of days total/partial, rounded up to the next integer.
day 1 $110
day 2 $110+60=$170
so S(4.5)=S(5)
S(4.5)=$110+$240=$350